OpenStudy (anonymous):

Find the exact solution for (√(5) - 1)/(x) = (√(5))/(2) Then, find the approximate solution.

OpenStudy (anonymous):

\[\frac{ \sqrt{5} - 1 }{ x } =\frac{ \sqrt{5} }{ 2 }\]

OpenStudy (anonymous):

reciprocal will also be equal to each other: \[\frac{x}{\sqrt5-1}=\frac{2}{\sqrt5}\] at this point multiply both sides by (sqrt5 -1)

OpenStudy (anonymous):

To be honest I don't know anything about this. Only very little about square roots...

OpenStudy (anonymous):

since it asks for an 'exact solution, you can keep the squareroots as they are. just have to get x by itself

OpenStudy (anonymous):

By multiplying /5 -1 on each side? How do I go about that?

OpenStudy (anonymous):

\[\cancel{(\sqrt5-1)}\times\frac{x}{\cancel{\sqrt5-1}}=\frac{2}{\sqrt5}\times (\sqrt5-1)\]

OpenStudy (anonymous):

it's exactly like it sounds. by multiplying each side by the same quantity you preserve the equal sign, then one side simplifies. it can also be considered a move across the equal sign, and above the division sign

OpenStudy (anonymous):

So from \[x = \frac{ 2 }{ \sqrt{5} } x (\sqrt{5} - 1)\] I just multiply that side and get the answer?

OpenStudy (anonymous):

Well one of the answers.